OpenStudy (anonymous):

What is the slope of the line passing through (-3, 5) and (5, -3)?

OpenStudy (anonymous):

slope = -1 slope = 0 slope = 1 no slope

OpenStudy (anonymous):

is there a graph?

OpenStudy (anonymous):

\[slope=\frac{ y_2-y_1 }{ x_2-x_1 }\]

OpenStudy (anonymous):

hold on

OpenStudy (anonymous):

i did the slope formula i didnt give me any of the answers that are the multiple choices

OpenStudy (anonymous):

-1 sorry

OpenStudy (anonymous):

its ok

OpenStudy (anonymous):

\[slope=\frac{ -3-5 }{ 5-(-3) }\]

OpenStudy (anonymous):

\[slope=\frac{ -8 }{ 8 }\]

OpenStudy (anonymous):

-8/8

OpenStudy (anonymous):

(-3, 5) and (5, -3) slope = (-3-5)/(5+3)=-8/8=-1

OpenStudy (anonymous):

-1

OpenStudy (anonymous):

yep :)

OpenStudy (anonymous):

ok thanks all

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

np

OpenStudy (anonymous):

no problem